Determine the maximum​ 30-year fixed-rate mortgage amount for which a couple could qualify if the rate is 4.37 percent. Assume they have other debt payments totaling $325 per month and a combined annual income of ​$70,000. Monthly escrow payments for real estate taxes and​ homeowner's insurance are estimated to be ​$162.​(Assume a 36 percent maximum of annual income for total debt and escrow​ payments.)
Answer #1

Maximum monthly income that can go towards debt = $70,000 / 12 * 0.36
= $2,100.

Total payments per month = $325 + $162 = $487

Monthly payment on the loan that they can afford = $2,100 - $487 = $1,613.

Nper = 30 * 12 = 360
Rate = 4.37% / 12
PMT = 1613
FV = 0

The maximum fixed rate mortgage can be calculated by using the following excel formula:
=PV(rate,nper,pmt,fv)
=PV(4.37%/12,360,-1613,0)
= $323,252.72

The maximum fixed rate mortgage = $323,252.72